solve the system bny substitution
x+y=12
y=2x

Answer by funmath(2933)

solve the system bny substitution
x+y=12
y=2x
Let y = 2x in the first equation and solve for x:
x+(2x)=12
3x=12
3x/3=12/3
x=4
Now, let x = 4 in the second equation and you'll find y:
y=2(4)
y=8
The solution to your system is: (x,y)=(4,8)
:
If you want to check, let x=4 and y=8 in both equations and see if the equations balance:
4+8=12 True
8=2(4) also true, we're right!!!
Happy Calculating!!!!